An electron of charge e is moving at a constant velocity u, along the x-axis. It enters a region of constant electric field E, which is pointing perpendicular to the x-axis. The electron moves in a parabola. Which of the following represents the focal length of the parabola? Neglect any effects due to gravity. (for a parabola of type x2=4ay, a is the focal length)


1
Expert's answer
2022-01-31T07:34:48-0500

The force acting on the electron:


"F=eE."

As the electron enters the region of the electric field, the force causes acceleration:


"a=F\/m=eE\/m."


Height above the sheet can be found as


"y=\dfrac12at^2."


Horizontal range can be found as


"x=ut\u2192t=x\/u."


Hence:


"y=\dfrac12ax^2/u^2=\dfrac{ax^2}{2u^2}=\dfrac{eEx^2}{2mu^2},\\\space\\
x^2=\dfrac{2mu^2}{eE}y=\dfrac{2mu^2}{eE}y=4·\dfrac{mu^2}{2eE}·y,"

the focal length is


"f=\dfrac{mu^2}{2eE}."